OpenStudy (anonymous):

csc t/sec^2t*sin t-sin t cost^2t/cos^2t

myininaya (myininaya):

what are we doing?

myininaya (myininaya):

simplifying?

OpenStudy (anonymous):

oh sorry. yes simplifying

OpenStudy (anonymous):

Put everything in terms of sin(t) and cos(t)...I can't quite identify if you're multiplying the denominator of the first term by sin(t) or the whole fraction. Could you please clarify with the equation editor as a response?

myininaya (myininaya):

yes i agree with modulus. i cant read.

myininaya (myininaya):

okay cool so cscx=1/sinx and 1/(secx)^2=(cosx)^2

myininaya (myininaya):

We have now that [(cost)^2/sint]*[(sint-sint(cost)^2/(cost)^2]

myininaya (myininaya):

we see that (cost)^2 cancel giving us (sint-sint(cost)^2)/sint

myininaya (myininaya):

But oh mine there is a sint in each term so we can cancel a sint from each term because sint/sint=1: (1-(cost)^2)/1

myininaya (myininaya):

But (sint)^2+(cost)^2=1 so 1-(cost)^2=(sint)^2

OpenStudy (anonymous):

thanks